The following data represents the age in weeks at which babies start crawling based on a sample of 12th parents. 30 2435 39 26 52 30 47 a. Calculate the average. b. If the sample standard deviation is s 9.67, the critical value is 1.65 and if the distribution is normally distributed calculate the 90% confidence interval.
Answer #1

Solution:

Sample average can be calculated as

=(52+47+30+37+44+56+35+26+39+39+26+28)/12

Average = 459/12 = 38.25

Solution(b)

Confidence interval can be calculated as

Xbar +/- zalpha/2 *S/sqrt(n)

38.25 +/- 1.65*9.67/sqrt(12)

38.25 +/- 4.61141

So 90% confidence interval is

33.6385 to 42.861